OpenStudy (fjr2099):

An acute angle θ is in a right triangle with sin θ = six sevenths. What is the value of cot θ?

OpenStudy (visarika23):

Sin(theta)=6/7=perpendicular/hypotenuse Base=sqr root(h^2-p^2) Cot(theta)=base/perpendicular